OpenStudy (anonymous):

Evaluate the expression below when x = 9. 3x + 2x2 - 9

OpenStudy (anonymous):

it depends. if its 3*9+2*9*2-9 or 3*9+2*2-9. Either way just replace x in the problem with 9 and solve.

OpenStudy (anonymous):

3x = 2x^2 - 9

OpenStudy (anonymous):

alright so then calculate each side. 3*9=? and 2*9^2-9=? if the ? are the same than it checks.

OpenStudy (anonymous):

i did it but idk if i did it right

OpenStudy (anonymous):

3*9=27, 9^2=81 so 2*81-9=153

OpenStudy (anonymous):

I second that ^answer^

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

but thats not a option

OpenStudy (anonymous):

Is it x^2 or just 2x2?

OpenStudy (anonymous):

If it is x^2 then I get 180 but if it is 2x2 I get 54. either one of those possible answers?

OpenStudy (anonymous):

they are both an option

OpenStudy (anonymous):

So does the equation of x raised to the second power or simply times 2? I dont know unless I see the original paper that the problem was printed on

OpenStudy (anonymous):

It's online

OpenStudy (anonymous):

Evaluate the expression below when x = 9. 3x + 2x2 - 9